# GOUHFI 2.0: A Next-Generation Toolbox for Brain Segmentation and Cortex Parcellation at Ultra-High Field MRI

### v1

@@ 1 Introduction
Historically, the only option that offered both whole-brain segmentation and cortical parcellation was FreeSurfer for T1w images at 1.5 and 3T. With the increased availability and affordability of GPU hardware, several neuroscientists have replaced FreeSurfer by its DL-equivalent FastSurfer, reducing by 25 times (or even more) computation time [undefaaf]. Additionally, since sub-millimeter resolution has become increasingly available, a more recent variation of FastSurfer, named FastSurferVINN, was proposed. FastSurferVINN segments T1w images acquired at 1.5 or 3T at native resolution [undefaag]. Both FreeSurfer and FastSurfer propose to segment the whole-brain into 35 subcortical regions as originally defined in [undefaz] and the cortex into 62 regions (31 for each hemisphere) following the DKT atlas [undefaan].
@@
As a result, over the past two decades, FreeSurfer, and more recently FastSurferVINN, have become the gold standard for automatic brain segmentation and cortical parcellation for T1w images at 1.5-3T and have been extensively used and validated by the neuroimaging community [undefaw, undefau, undefaam, undefaaae].
@@
Another important neuroimaging toolbox is the Advanced Normalization Tools (ANTs) which has been broadly utilized to study and analyze brain MR images [undefar]. Lately, a novel implementation of the toolbox employing Deep Learning (DL) has been made available under the names of ANTsRNet/ANTsPyNet corresponding to R and Python implementations of the tool respectively [undefaaaa]. While the automatic brain segmentation available through these tools segments the brain into six tissue classes (Cerebrospinal Fluid (CSF), Gray Matter/Cortex (GM), White Matter (WM), deep GM, brainstem and cerebellum) compared to 35 for FastSurferVINN, the tool antspynet.desikan_killiany_tourville_labeling enables the parcellation of the cortex under the DKT convention, mimicking the FastSurferVINN ecosystem.
@@
More recently, a novel automatic brain segmentation technique, called SynthSeg+ (also available in the FreeSurfer ecosystem), has been proposed where a novel DL approach called domain randomization (DR) is used to train a model using synthetic images [undefat]. Following the same labeling convention as defined by FreeSurfer, SynthSeg+ is able to segment the brain and cortex in a few minutes like FastSurfer. The main difference with FastSurfer resides in SynthSeg+’s ability to segment images of any contrast or resolution with a fixed output resolution of 1.0 mm3, allowing the use of non-T1w contrasts or lower resolution images as input. SynthSeg+ allows neuroscientists with clinical images acquired with lower image quality than usually acquired in research environments to be used for quantitative analyses like volumetry, which have been impossible previously. Even if proposed recently, the DR approach applied to brain segmentation has been used by several segmentation tools, showing promising avenues for the development of novel techniques for a varied range of applications [undefaaab, undefaad, undefaaa].
@@
While the previously mentioned techniques have been successfully applied and validated, none of them have been widely applied to UHF-MRI (i.e., $\geq$7T). UHF-MRI has been increasingly available in the last decade and used for large neuroimaging studies like the Human Connectome Project (HCP) [undefaaac]. Despite its several advantages, like higher SNR, contrast and spatial resolution, UHF images typically suffer from significant transmit radiofrequency (RF) inhomogeneities compared to lower fields [undefaaz]. This results in significant signal and contrast inhomogeneities observed across the image. Although recent developments in parallel transmit (pTx) RF pulses have substantially improved this issue [undefaae], their use remains limited and yet to be applied to large neuroimaging studies.
@@
Considering this lack of automatic segmentation methods developed for UHF-MRI, we have recently proposed GOUHFI: Generalized and Optimized segmentation tool for Ultra-High Field Images [undefaaa]. By integrating and refining the DR strategy from SynthSeg+ with datasets more tailored to the UHF context for training and testing, GOUHFI proposes an inhomogeneity-resistant, contrast- and resolution-agnostic brain segmentation tool capable of segmenting the brain into 35 labels following the FreeSurfer label convention. Considering there is yet to be a standard high-resolution at UHF and many variations of T1w protocols are used in the community, the contrast and resolution agnosticity of GOUHFI revealed itself as a promising tool for segmenting UHF images, especially in cases with significant signal inhomogeneities. GOUHFI is available online as an open-source toolbox at the following address: https://github.com/mafortin/GOUHFI.
@@
While extensively validated on six contrasts and seven resolutions at 3T, 7T and 9.4T, GOUHFI was mainly validated on healthy and young adults. However, the STRAT-PARK dataset [undefaax], which comprises patients with Parkinson’s disease (PDP) and age-matched healthy controls (HC), was used to evaluate GOUHFI in a clinical cohort setting. While GOUHFI performed well in most subjects, few exposed GOUHFI’s limitations. More precisely, subjects with substantially enlarged lateral ventricles (also frequently observed for healthy but older brain anatomies), were challenging for GOUHFI. Indeed, proper delineation of the caudate, thalamus and hippocampus adjacent to large ventricles was sub-optimal, resulting in poor identification of the actual borders between the neighboring regions. While impacting a minority of test cases, this observation raised the need for a more exhaustive validation on non-healthy and older brain anatomies for GOUHFI.
@@ 2.1 Datasets
Most of the data used in this study are the same as in [undefaaa]. An overview of all these datasets is provided in Table 1 below, where the training and test datasets are described. In this study, three datasets were added, one for training and two for testing, and are presented in the following subsections.
Table 1: Summary of the datasets used for training and testing in this work. The new datasets are in bold. The table lists the field strength, resolution, contrast, subject type, vendor, and number of subjects. ASD: Autism Spectrum Disorder, AD: Alzheimer’s Disease, PDP: Parkinson’s Disease Patients, SCA: SpinoCerebellar Ataxia.
@@
STRAT-PARK 7T (1Tx) (0.75 mm)3 T1w PDP/Healthy Siemens 45
@@ 2.1.1 Open Access Series of Imaging Studies: OASIS2 & OASIS1-Test-Retest
A subset of 32 subjects from the Open Access Series of Imaging Studies 2 (OASIS2) [undefaas] was selected in order to increase the anatomical variation in the training dataset by adding demented and non-demented subjects aged between 72 and 97 years old. For each subject, a 1.0 mm3 T1w MPRAGE was acquired with a 1.5T Vision scanner (Siemens Healthineers, Erlangen, Germany). Subjects were selected based on the presence of substantially enlarged ventricles, creating anatomical variations impossible to create through the generation of synthetic images with the generative model. The complete OASIS2 dataset and information about the dataset can be accessed online [^1] https://sites.wustl.edu/oasisbrains/home/oasis-2/.
A second subset of the OASIS data was utilized for testing GOUHFI 2.0. The manually delineated cortical parcellation done on 20 subjects from the OASIS1-Test-Retest (OASIS-TRT-20) dataset, also known as the MindBoggle-101 dataset [undefaan], was used. The manual delineations were made from a 1.0 mm3 T1w MPRAGE and follow the DKT protocol (31 cortical labels per hemisphere), which is the same protocol used by FastSurferVINN. To the best of our knowledge, this is the only set of manually segmented cortex parcellations following this protocol available online [^2] https://mindboggle.info/data.html.
@@ 2.1.2 SCAIFIELD-NPC: Spinocerebellar Ataxia - Norwegian Patient Cohort
In this study, we used the SCAIFIELD Norwegian Patient Cohort (SCAIFIELD-NPC) as a test dataset. While the imaging protocol is exactly the same as for the SCAIFIELD protocol, the difference resides in the fact that it only includes three SCA subjects instead of healthy controls. This dataset was used to qualitatively assess the performance of FastSurferVINN, SynthSeg+ and GOUHFI 2.0 in a clinical cohort acquired at 7T.
@@ 2.2 Data processing
Except for the modifications mentioned below, the rest of the processing of the training data was kept untouched from that described in the original GOUHFI publication and refer the reader to it [undefaaa].
@@ 2.2.1 Initial training label maps creation with FastSurferVINN
All T1w images used in this study were segmented using FastSurferVINN [undefaag] (v2.3.0) with the –seg_only flag in order to produce automatic whole brain segmentations into 95 labels (33 subcortical labels and 62 cortex labels following the DKT labeling protocol). The list of labels produced by FastSurferVINN, and used in this work, is available in Table 5 of Appendix A of [undefaag].
@@ 2.2.3 Training datasets: Brain segmentation and cortex parcellation
In contrast with its original implementation, GOUHFI 2.0 now utilizes two 3D U-Nets [undefaav] where one network is used for the brain segmentation task and the other one for the cortex parcellation. Regarding the first network (A in Figure 2), the training label maps were used to generate synthetic images following the domain randomization (DR) approach proposed in [undefas] adapted to the UHF-MRI context as explained in [undefaaa]. While the training corpus consisted of synthetic images, the validation set was composed of the corresponding real masked T1w image for each subject, except for the HCP-YA and SCAIFIELD, where the coregistered T2w and PDw images were used. Validating the model on real MR images allows to check if the model is able to learn domain-agnostic features that translates well into the actual intended usage with real MR images and avoids learning features specific to synthetic images only.
Moreover, [undefaag] showed that enlarging the training corpus was the key factor for improving both performance and generalizability of DL segmentation models. Motivated by this, we designed two training dataset variants for the brain segmentation task in GOUHFI 2.0. The first variant, GOUHFI 2.0-n1, followed the original GOUHFI strategy, where one synthetic image was generated per input label map for the generative model (238 synthetic images). Since the generation procedure synthesizes images with widely varying contrasts, artifacts, and anatomical shapes, a second variant where two synthetic images per input label map was tested. This second variant, named GOUHFI 2.0-n2, had a total of 476 synthetic images in its training corpus.
@@ 2.4.1 Competing methods
For brain segmentation, FastSurferVINN (v2.3.0) [undefaag], SynthSeg+ [undefat] and the original GOUHFI were used for qualitative comparison with the two variants of GOUHFI 2.0. For the quantitative comparison, FastSurferVINN was excluded due to its limitations at UHF and inability to segment non-T1w contrasts. While all above-mentioned methods segment at native input image resolution, SynthSeg+ generates segmentation outputs at a standardized resolution of 1.0 mm3, independent of the input resolution. Consequently, to obtain label maps at native resolution and allow quantitative comparisons, the same external up-sampling strategy using one-hot encoded 3D linear interpolation for label maps as done for GOUHFI was implemented in-house for SynthSeg+’s label maps. For cortex parcellation, the mentioned methods were used, in addition to ANTsPyNet [undefaaaa].
@@ 2.4.2 Quantitative evaluation
In order to assess the quality of the segmentations produced by GOUHFI 2.0, the Dice-Sørensen Similarity Coefficient (DSC) [undefax, undefaaw], which measures the overlap between two segments (with a value of 1 being a perfect overlap between the two segments) was computed with the following equation:
@@
Moreover, the Average Surface Distance (ASD) [undefaau], where a value of 0 represents a perfect alignment of both surfaces evaluated, was computed with the following equation:
@@ 2.4.3 Total intracranial volume estimation and volumetry analysis
The volume of individual brain structures was computed by summing all voxels present in the final label maps and multiplying it by the voxel size. For the total intracranial volume (TIV), all structures were summed together. In order to assess the TIV estimation from both GOUHFI 2.0 variants, the computed TIV values were compared with the estimation from SynthSeg+ and SPM12 [undefaq], the latter being considered the gold standard technique [undefaaq].
As part of the assessment of modifying the training corpus for GOUHFI 2.0, the same volumetric study as performed for the original GOUHFI was done on the STRAT-PARK clinical cohort with PDP at 7T. For both HC and PDP, the mean group volume for the putamen, amygdala and hippocampus, normalized by TIV, were computed based on the segmentations produced by FastSurferVINN (reference), GOUHFI 2.0 and SynthSeg+. These ROIs were selected based on the literature for PD [undefaal, undefaat, undefaac]. A Mann-Whitney U test [undefaar] with Bonferroni-corrected p-values [undefav] was computed to measure the statistical differences between each group.
@@ 3 Results
Figure 4 shows the performance of all segmentation techniques tested for a PDP subject acquired at 1 Tx 7T MP2RAGE. In cases where subjects would exhibit enlarged ventricles either due to pathology or age, GOUHFI revealed two main limitations unique to it. First, as shown in the blue arrows in Figure 4, GOUHFI segmented the caudate and some parts of the thalamus inside the lateral ventricles for one subject. It is important to mention that this behavior was only observed in one subject in the STRAT-PARK dataset (out of 45). Moreover, the other limitation of GOUHFI was its frequent, although not systematic, poor delineation of the boundary between the hippocampus and inferior lateral ventricles (red/yellow/green arrows on Figure 4). For both variants of GOUHFI 2.0, none of these limitations were observed. However, a slightly improved hippocampus segmentation and inhomogeneity resistance (orange arrow) was observed for GOUHFI 2.0-n2 compared to GOUHFI 2.0-n1.
Figure 4: Figure presenting the differences in segmentations between FastSurferVINN, SynthSeg+, GOUHFI, GOUHFI 2.0-n1 and GOUHFI 2.0-n2, computed from an N4-corrected MP2RAGE image highly affected by signal inhomogeneities. The example is from a PDP subject in the STRAT-PARK dataset, exhibiting substantial ventricular enlargement. Red, yellow and green arrows point to the quality of the hippocampus-inferior lateral ventricle delineations across techniques. Blue arrows demonstrate the repeated improvement of the segmentation of enlarged ventricles throughout GOUHFI’s versions. Orange arrows demonstrate the performance of the different algorithms regarding the segmentation of the cortex in presence of highly inhomogeneous signal.

### v2

@@ 1 Introduction
Historically, the only option that offered both whole-brain segmentation and cortical parcellation was FreeSurfer for T1w images at 1.5 and 3T. With the increased availability and affordability of GPU hardware, several neuroscientists have replaced FreeSurfer by its DL-equivalent FastSurfer, reducing by 25 times (or even more) computation time [18]. Additionally, since sub-millimeter resolution has become increasingly available, a more recent variation of FastSurfer, named FastSurferVINN, was proposed. FastSurferVINN segments T1w images acquired at 1.5 or 3T at native resolution [19]. Both FreeSurfer and FastSurfer propose to segment the whole-brain into 35 subcortical regions as originally defined in [12] and the cortex into 62 regions (31 for each hemisphere) following the DKT atlas [26].
@@
As a result, over the past two decades, FreeSurfer, and more recently FastSurferVINN, have become the gold standard for automatic brain segmentation and cortical parcellation for T1w images at 1.5-3T and have been extensively used and validated by the neuroimaging community [8, 6, 25, 46].
@@
Another important neuroimaging toolbox is the Advanced Normalization Tools (ANTs) which has been broadly utilized to study and analyze brain MR images [3]. Lately, a novel implementation of the toolbox employing Deep Learning (DL) has been made available under the names of ANTsRNet/ANTsPyNet corresponding to R and Python implementations of the tool respectively [41]. While the automatic brain segmentation available through these tools segments the brain into six tissue classes (Cerebrospinal Fluid (CSF), Gray Matter/Cortex (GM), White Matter (WM), deep GM, brainstem and cerebellum) compared to 35 for FastSurferVINN, the tool antspynet.desikan_killiany_tourville_labeling enables the parcellation of the cortex under the DKT convention, mimicking the FastSurferVINN ecosystem.
@@
More recently, a novel automatic brain segmentation technique, called SynthSeg+ (also available in the FreeSurfer ecosystem), has been proposed where a novel DL approach called domain randomization (DR) is used to train a model using synthetic images [5]. Following the same labeling convention as defined by FreeSurfer, SynthSeg+ is able to segment the brain and cortex in a few minutes like FastSurfer. The main difference with FastSurfer resides in SynthSeg+’s ability to segment images of any contrast or resolution with a fixed output resolution of 1.0 mm3, allowing the use of non-T1w contrasts or lower resolution images as input. SynthSeg+ allows neuroscientists with clinical images acquired with lower image quality than usually acquired in research environments to be used for quantitative analyses like volumetry, which have been impossible previously. Even if proposed recently, the DR approach applied to brain segmentation has been used by several segmentation tools, showing promising avenues for the development of novel techniques for a varied range of applications [42, 16, 13].
@@
While the previously mentioned techniques have been successfully applied and validated, none of them have been widely applied to UHF-MRI (i.e., $\geq$7T). UHF-MRI has been increasingly available in the last decade and used for large neuroimaging studies like the Human Connectome Project (HCP) [43]. Despite its several advantages, like higher SNR, contrast and spatial resolution, UHF images typically suffer from significant transmit radiofrequency (RF) inhomogeneities compared to lower fields [40]. This results in significant signal and contrast inhomogeneities observed across the image. Although recent developments in parallel transmit (pTx) RF pulses have substantially improved this issue [17], their use remains limited and yet to be applied to large neuroimaging studies.
@@
Considering this lack of automatic segmentation methods developed for UHF-MRI, we have recently proposed GOUHFI: Generalized and Optimized segmentation tool for Ultra-High Field Images [13]. By integrating and refining the DR strategy from SynthSeg+ with datasets more tailored to the UHF context for training and testing, GOUHFI proposes an inhomogeneity-resistant, contrast- and resolution-agnostic brain segmentation tool capable of segmenting the brain into 35 labels following the FreeSurfer label convention. Considering there is yet to be a standard high-resolution at UHF and many variations of T1w protocols are used in the community, the contrast and resolution agnosticity of GOUHFI revealed itself as a promising tool for segmenting UHF images, especially in cases with significant signal inhomogeneities. GOUHFI is available online as an open-source toolbox at the following address: https://github.com/mafortin/GOUHFI.
@@
While extensively validated on six contrasts and seven resolutions at 3T, 7T and 9.4T, GOUHFI was mainly validated on healthy and young adults. However, the STRAT-PARK dataset [38], which comprises people with Parkinson’s disease (PwP) and age-matched healthy controls (HC), was used to evaluate GOUHFI in a clinical cohort setting. While GOUHFI performed well in most subjects, few exposed GOUHFI’s limitations. More precisely, subjects with substantially enlarged lateral ventricles (also frequently observed for healthy but older brain anatomies), were challenging for GOUHFI. Indeed, proper delineation of the caudate, thalamus and hippocampus adjacent to large ventricles was sub-optimal, resulting in poor identification of the actual borders between the neighboring regions. While impacting a minority of test cases, this observation raised the need for a more exhaustive validation on non-healthy and older brain anatomies for GOUHFI.
@@ 2.1 Datasets
Most of the data used in this study are the same as in [13]. An overview of all these datasets is provided in Table 1 below, where the training and test datasets are described. In this study, three datasets were added, one for training and two for testing, and are presented in the following subsections.
Table 1: Summary of the datasets used for training and testing in this work. The new datasets are in bold. The table lists the field strength, resolution, contrast, subject type, vendor, and number of subjects. ASD: Autism Spectrum Disorder, AD: Alzheimer’s Disease, PwP: People with Parkinson’s Disease, SCA: SpinoCerebellar Ataxia.
@@
STRAT-PARK 7T (1Tx) (0.75 mm)3 T1w PwP/Healthy Siemens 45
@@ 2.1.1 Open Access Series of Imaging Studies: OASIS2 & OASIS1-Test-Retest
A subset of 32 subjects from the Open Access Series of Imaging Studies 2 (OASIS2) [32] was selected in order to increase the anatomical variation in the training dataset by adding demented and non-demented subjects aged between 72 and 97 years old. For each subject, a 1.0 mm3 T1w MPRAGE was acquired with a 1.5T Vision scanner (Siemens Healthineers, Erlangen, Germany). Subjects were selected based on the presence of substantially enlarged ventricles, creating anatomical variations impossible to create through the generation of synthetic images with the generative model. The complete OASIS2 dataset and information about the dataset can be accessed online [^1] https://sites.wustl.edu/oasisbrains/home/oasis-2/.
A second subset of the OASIS data was utilized for testing GOUHFI 2.0. The manually delineated cortical parcellation done on 20 subjects from the OASIS1-Test-Retest (OASIS-TRT-20) dataset, also known as the MindBoggle-101 dataset [26], was used. The manual delineations were made from a 1.0 mm3 T1w MPRAGE and follow the DKT protocol (31 cortical labels per hemisphere), which is the same protocol used by FastSurferVINN. To the best of our knowledge, this is the only set of manually segmented cortex parcellations following this protocol available online [^2] https://mindboggle.info/data.html.
@@ 2.1.2 SCAIFIELD-NPC: Spinocerebellar Ataxia - Norwegian Patient Cohort
In this study, we used the SCAIFIELD Norwegian Patient Cohort (SCAIFIELD-NPC) as a test dataset. While the imaging protocol is exactly the same as for the SCAIFIELD protocol, the difference resides in the fact that it only includes three SCA subjects instead of healthy controls. SCA is a heterogeneous group of neurodegenerative genetic disorders characterized by progressive cerebellar and brain atrophy [27]. Because SCA induces substantial brain atrophy, this dataset was used strictly for a qualitative assessment of the segmentation outputs from FastSurferVINN, SynthSeg+, and GOUHFI2.0, rather than for quantitative performance comparison, in a 7T clinical cohort.
@@ 2.2 Data processing
Except for the modifications mentioned below, the rest of the processing of the training data was kept untouched from that described in the original GOUHFI publication and refer the reader to it [13].
@@ 2.2.1 Initial training label maps creation with FastSurferVINN
All T1w images used in this study were segmented using FastSurferVINN [19] (v2.3.0) with the –seg_only flag in order to produce automatic whole brain segmentations into 95 labels (33 subcortical labels and 62 cortex labels following the DKT labeling protocol). The list of labels produced by FastSurferVINN, and used in this work, is available in Table 5 of Appendix A of [19].
@@ 2.2.3 Training datasets: Brain segmentation and cortex parcellation
In contrast with its original implementation, GOUHFI 2.0 now utilizes two 3D U-Nets [35] where one network is used for the brain segmentation task and the other one for the cortex parcellation. Regarding the first network (A in Figure 2), the training label maps were used to generate synthetic images following the domain randomization (DR) approach proposed in [4] adapted to the UHF-MRI context as explained in [13]. While the training corpus consisted of synthetic images, the validation set was composed of the corresponding real masked T1w image for each subject, except for the HCP-YA and SCAIFIELD, where the coregistered T2w and PDw images were used. Validating the model on real MR images allows to check if the model is able to learn domain-agnostic features that translates well into the actual intended usage with real MR images and avoids learning features specific to synthetic images only.
Moreover, [19] showed that enlarging the training corpus was the key factor for improving both performance and generalizability of DL segmentation models. Motivated by this, we designed two training dataset variants for the brain segmentation task in GOUHFI 2.0. The first variant, GOUHFI 2.0-n1, followed the original GOUHFI strategy, where one synthetic image was generated per input label map for the generative model (238 synthetic images). Since the generation procedure synthesizes images with widely varying contrasts, artifacts, and anatomical shapes, a second variant where two synthetic images per input label map was tested. This second variant, named GOUHFI 2.0-n2, had a total of 476 synthetic images in its training corpus.
@@ 2.4.1 Competing methods
For brain segmentation, FastSurferVINN (v2.3.0) [19], SynthSeg+ [5] and the original GOUHFI were used for qualitative comparison with the two variants of GOUHFI 2.0. For the quantitative comparison, FastSurferVINN was excluded due to its limitations at UHF and inability to segment non-T1w contrasts. While all above-mentioned methods segment at native input image resolution, SynthSeg+ generates segmentation outputs at a standardized resolution of 1.0 mm3, independent of the input resolution. Consequently, to obtain label maps at native resolution and allow quantitative comparisons, the same external up-sampling strategy using one-hot encoded 3D linear interpolation for label maps as done for GOUHFI was implemented in-house for SynthSeg+’s label maps. For cortex parcellation, the mentioned methods were used, in addition to ANTsPyNet [41].
@@ 2.4.2 Quantitative evaluation
In order to assess the quality of the segmentations produced by GOUHFI 2.0, the Dice-Sørensen Similarity Coefficient (DSC) [9, 37], which measures the overlap between two segments (with a value of 1 being a perfect overlap between the two segments) was computed with the following equation:
@@
Moreover, the Average Surface Distance (ASD) [34], where a value of 0 represents a perfect alignment of both surfaces evaluated, was computed with the following equation:
@@ 2.4.3 Total intracranial volume estimation and volumetry analysis
The volume of individual brain structures was computed by summing all voxels present in the final label maps and multiplying it by the voxel size. For the total intracranial volume (TIV), all structures were summed together. In order to assess the TIV estimation from both GOUHFI 2.0 variants, the computed TIV values were compared with the estimation from SynthSeg+ and SPM12 [2], the latter being considered the gold standard technique [30].
As part of the assessment of modifying the training corpus for GOUHFI 2.0, the same volumetric study as performed for the original GOUHFI was done on the STRAT-PARK clinical cohort with PwP at 7T. For both HC and PwP, the mean group volume for the putamen, amygdala and hippocampus, normalized by TIV, were computed based on the segmentations produced by FastSurferVINN (reference), GOUHFI 2.0 and SynthSeg+. These ROIs were selected based on the literature for PD [24, 33, 15]. A Mann-Whitney U test [31] with Bonferroni-corrected p-values [7] was computed to measure the statistical differences between each group.
@@ 3 Results
Figure 4 shows the performance of all segmentation techniques tested for a PwP subject acquired at 1 Tx 7T MP2RAGE. In cases where subjects would exhibit enlarged ventricles either due to pathology or age, GOUHFI showed two main limitations which were specific to this approach. First, in a single subject of the 45 in the STRAT-PARK dataset, GOUHFI segmented the caudate and some parts of the thalamus inside the lateral ventricles (blue arrows in Figure 4). It is important to mention that this behavior was only observed in one subject in the STRAT-PARK dataset (out of 45). Second, GOUHFI often resulted in poor delineation of the boundary between the hippocampus and inferior lateral ventricles (red/yellow/green arrows on Figure 4). None of the two variants of GOUHFI 2.0 exhibited these limitations. However, a slightly improved hippocampus segmentation and inhomogeneity resistance (orange arrow) was observed for GOUHFI 2.0-n2 compared to GOUHFI 2.0-n1.
Figure 4: Figure presenting the differences in segmentations between FastSurferVINN, SynthSeg+, GOUHFI, GOUHFI 2.0-n1 and GOUHFI 2.0-n2, computed from an N4-corrected MP2RAGE image highly affected by signal inhomogeneities. The example is from a PwP subject in the STRAT-PARK dataset, exhibiting substantial ventricular enlargement. Red, yellow and green arrows point to the quality of the hippocampus-inferior lateral ventricle delineations across techniques. Blue arrows demonstrate the repeated improvement of the segmentation of enlarged ventricles throughout GOUHFI’s versions. Orange arrows demonstrate the performance of the different algorithms regarding the segmentation of the cortex in presence of highly inhomogeneous signal.